OpenStudy (anonymous):

How many ordered triples $(x,y,z)$ of integers are there such that $\sqrt{x^2 + y^2 + z^2} = 7$? Does the question have a geometric interpretation? Please help! Thanks!

OpenStudy (jchick):

First square both sides.

OpenStudy (jchick):

x^2 + y^2 + z^2 =7^2

OpenStudy (jchick):

then we will end up with a sphere with radius 7

OpenStudy (anonymous):

so there are many triples. Is it correct?

OpenStudy (jchick):

You will end up with (0,0,7) ( 0, 7, 0) and ( 7, 0 , 0 ) as solutions

OpenStudy (jchick):

( 0, 0 , -7 ) ( 0,0,7) (0,7 , 0) ( 0 , -7 , 0 ) (7,0,0) , (-7, 0, 0)

OpenStudy (jchick):

Two of the principal solutions are 0,0,7 and 6,3,2 , then you just rearrange them and then you take their negatives.

OpenStudy (jchick):

This leaves you with 54 integer solutions

OpenStudy (anonymous):

I got it. Thank you!!

OpenStudy (jchick):

No problem!

OpenStudy (kainui):

@jchick how do you know there aren't more?

OpenStudy (jchick):

we know that (6,3,2) is a solution, and we can rearrange it , since they are symmetric x^2 + y^2 + z^2 = 7^2 6^2 + 3^2 + 2^2 = 7^2 3^2 + 6^2 + 2^2 = 7^2

OpenStudy (jchick):

each rearrangement of 6,3,2 is another solution x = 6 y = 3 z = 2 x = 6 , y = 2 , z = 3

OpenStudy (jchick):

6,3,2 6,2,3 3, 6,2 3,2,6 2,6,3 2,3,6

OpenStudy (jchick):

(0,0,7) , there are three ways to rerarrange that. (0,7,0) (7,0,0) , and times 2 since there is one neg/pos each so 3!*2*2*2 + 3 * 2 is the number of integer solutions

OpenStudy (jchick):

@Kainui